OpenStudy (twizttiez):

A hard candy has a radius of 9 mm. The candy has a spherical soft center with a radius of 6 mm, with a hard shell surrounding it. What is the volume of the hard shell? Use 3.14 to approximate pi and express your final answer in hundredths. ??? mm^3

OpenStudy (anonymous):

So just do total volume of hard candy - volume of soft center

OpenStudy (twizttiez):

Sorry but like I'm not good at math :'( could you help me with the formula??

OpenStudy (anonymous):

Volume of a sphere is \[\frac{ 4 }{ 3 }\pi*r^3\] Just put in the radius value and calculate in a calculator.

OpenStudy (twizttiez):

Okay one second!

OpenStudy (twizttiez):

288??

OpenStudy (twizttiez):

288pi

OpenStudy (anonymous):

How in the world did you get that answer?

OpenStudy (twizttiez):

I used 6 as my radius :/

OpenStudy (twizttiez):

With 9 I got 972pi

OpenStudy (anonymous):

You should have done\[(\frac{ 4 }{ 3 }\pi*9^3)-(\frac{ 4 }{ 3 }\pi*6^3)\] And you are supposed to use 3.14 as pi, so pi shouldn't be in the answer itself.

OpenStudy (anonymous):

Ok, I see what you did wrong. Subtract 288pi from 972 pi and then multiply the answer by 3.14

OpenStudy (twizttiez):

Ohhhhhh okay thanks so much

OpenStudy (anonymous):

What did you get for your answer?

OpenStudy (twizttiez):

Wait do I subtract 972 by 288 or the other way around?

OpenStudy (twizttiez):

???

OpenStudy (anonymous):

972-288

OpenStudy (twizttiez):

Okay thanks

OpenStudy (anonymous):

You're welcome

OpenStudy (twizttiez):

I got 2148.84937506

OpenStudy (twizttiez):

Right @LegendarySadist

OpenStudy (twizttiez):

6747.38703769 this is what I got after I multiplied it by pi